T/4+T/3=7/12
Solving Mutli-Step Equations.
The answer it T=1 correct?

Found 2 solutions by mananth, lagbens:
Answer by mananth(16946)

T=1 is correct

Answer by lagbens(14)

T/4+T/3=7/12.
let's solve for T.
we can multiply each term from left to right by 12.
we'd get 3T+4T=7(add the T terms up)
7T=7(divide both side by 7)
T=1 (final answer)